My fish tank is 1 foot wide 4 foot long and 21 inches tall how many gallons is my fish tank?
Answer by fcabanski(874)

Volume is length * width * height (depth).


In this case l = 4 ft., w = 1ft, h=21 inches. Convert inches to feet (divide by 12). So h = 21/12 = 1.75 ft.


Volume - 4*1*1.75= 7 cubic feet.


There are appox. 7.48 gallons for each cubic foot. So multiply:


7*7.48=52.36 gallons.

I provide online tutoring ($30/hr) and personal problem solving
($3.50-$5.50 per problem. Contact me (check the profile "website" which is
my email address.